q1 = 3.1 uc ,q2=11.4 uc ,q3= - 2.8 uc , d1= 4.9 cm ,

 d2=8.1 cm ,the net electric force on q1


1
Expert's answer
2021-04-19T17:44:30-0400

To be given in question

q1=3.1"\mu" "C"

q2=11.4"\mu" "C"

q3=-2.8"\mu" "C"

d2=4.9 cm ="4.9\times10^{-2} meter"

d2=8.1 cm=8.1"\times10^{-2} meter"

To be asked in question

"(F_{net}){q_{1}}=?"

We know that

Electric force between charge q1and q2

"F_{12} =k\frac{q_{1}q_{2}}{r_{12}^2}"

Put values

"F_{12}=\frac{9\times10^9\times3.1\times10^{-6}\times11.4\times10^{-6}}{{.049}^2}"



"F_{12}=132.46N"

Electric force between charge q1and q3

"F_{13}=\frac{ kq_{1}q_{3}}{r_{13}}"


Put value

"F_{13}=-\frac{9\times10^9\times3.1\times10^{-6}\times2.8\times10^{-6}}{{.081}^2}"

"F_{13}= -11.9067N"

Net force

"F=F_{12}+F_{13}"

"F_{net}=132.46-11.9067"

"F_{net}=120.55 N"